# Automated forest inventory: analysis of high-density airborne LiDAR point clouds with 3D deep learning

### 22 Dec 2023 v1

@@ Abstract
Detailed forest inventories are critical for sustainable and flexible management of forest resources, to conserve various ecosystem services. Modern airborne laser scanners deliver high-density point clouds with great potential for fine-scale forest inventory and analysis, but automatically partitioning those point clouds into meaningful entities like individual trees or tree components remains a challenge. The present study aims to fill this gap and introduces a deep learning framework that is able to perform such a segmentation across diverse forest types and geographic regions. From the segmented data, we then derive relevant biophysical parameters of individual trees as well as stands. The system has been tested on FOR-Instance, a dataset of point clouds that have been acquired in five different countries using surveying drones. The segmentation back-end achieves over 85% F-score for individual trees, respectively over 73% mean IoU across five semantic categories: ground, low vegetation, stems, live branches and dead branches. Building on the segmentation results our pipeline then densely calculates biophysical features of each individual tree (height, crown diameter, crown volume, DBH, and location) and properties per stand (digital terrain model and stand density). Especially crown-related features are in most cases retrieved with high accuracy, whereas the estimates for DBH and location are less reliable, due to the airborne scanning setup.
@@ 1 Introduction
Collecting and maintaining forest inventories was one of the first applications to adopt airborne laser scanning (ALS) as an operational tool, at local (Næsset et al., 2004), regional and national scales (Kangas et al., 2018). The two main approaches to extract forest inventory data from Light Detection and Ranging (LiDAR) point clouds are the area-based approach (ABA, Næsset, 2002) and individual tree detection (ITD, Hyyppa et al., 2001). ABA operates by estimating statistical indicators from the points within a given area (e.g., a grid cell). In contrast, ITD (also referred to as single tree inventory) relies on the detection and segmentation of individual trees in the point cloud (Persson et al., 2002). Several methods have been developed for ITD, whose accuracies differ depending on the forest conditions and the point density (Vauhkonen et al., 2011; Kaartinen et al., 2012). Most ITD algorithms operate on a rasterised canopy height model (CHM) derived from the highest ALS returns, representing the canopy top layer. Consequently, they are limited to the dominant canopy trees and tend to miss an important share of intermediate and understory trees. Despite the large body of research on the topic, and different data resolution requirements (Coomes et al., 2017) between ABA and ITC, the operational uptake of ITD inventories has so far been limited, due to (1) the bias introduced by the relatively low detection rates, caused by the omission of non-dominant trees, and (2) a lack of transferability across different forest conditions and datasets.
@@
Traditionally, low-resolution ALS-based inventories have been based on rather sparse point clouds (0.5 – 5 pts/m2) for ABA, and slightly denser for ITD (10 – 20 pts/m2). With the rapid development of sensor technology, very high density ALS (ALS-HD) has become available and offers new opportunities  (Kellner et al., 2019). In this context, is important to clarify that, while collected using airborne platforms such as drones, or helicopters  (Hyyppä et al., 2022; Persson et al., 2022), ALS-HD data differ compared to ALS data in the acquisition which is typically done at lower altitudes, with lower flight speed, and with larger lateral overlap between adjacent flight strips. These data are characterized by higher resolution point clouds compared to ALS data, with densities ranging from 500 up to 10,000 pts/m2 (e.g. Puliti et al. (2022a)). While available at a smaller geographical scale (e.g. stand level) compared to ALS, These data represent an emerging alternative data source for ITC inventories as they provide an unprecedented (for airborne data) level of detail and close the gap between large-area ALS and high-resolution terrestrial laser scanning (TLS). While this study relied on laser scanning data captured from unmanned aerial vehicles (UAVs), we referred to them as ALS-HD since the main difference compared to previous studies lies in the increased point cloud resolution rather than the airborne acquisition platform.
@@
The increased amount of 3D structural detail, with distinctively visible individual trees throughout the vertical canopy profile, has sparked renewed interest in the pursuit of single tree inventories, and possibly even completely airborne inventories based on dense retrievals of the variables of interest from the point clouds (Liang et al., 2019; Puliti et al., 2020). In addition to the measurement of common forestry variables like the diameter at breast height (DBH, Kuželka et al., 2020) and tree biomass (Brede et al., 2019), recent research has highlighted the potential to derive a suite of useful tree analytics and insights from per-tree ALS-HD data, such as species (Hakula et al., 2023), stem curve (Hyyppä et al., 2022), growth trajectory (Puliti et al., 2022b), and breeding-related traits (du Toit et al., 2023). The promise of high-resolution scans for ITD is evident, but there is a bottleneck: to measure the properties of each tree, one first needs a sufficiently accurate and transferable method to segment the raw scan data into individual trees.
@@
Parallel to the rise of high-density ALS, we have seen an explosion of interest in deep machine learning algorithms for processing 3D point clouds in general, and forest scans in particular (Table 1). Over the last decade, the superior capabilities of deep neural networks when it comes to representing and analysing structured data have brought about transformative changes across a range of science and engineering disciplines (Wang et al., 2023). Deep learning also holds great promise for tree segmentation, and indeed there has been a recent trend to adopt it for the ongoing efforts towards automated forest inventories (Xi et al., 2018; Chen et al., 2021b; Krisanski et al., 2021b; Sun et al., 2022; Chang et al., 2022; Wang and Bryson, 2023; Zhang et al., 2023; Kim et al., 2023; Jiang et al., 2023; Wielgosz et al., 2023; Straker et al., 2023).
@@
In the present work, we consolidate data-driven segmentation and structuring of forest point clouds into a single, integrated deep learning framework. We argue that a key ingredient for an efficient, automatic airborne forest inventory and management system is a comprehensive segmentation engine. In this way, reference data for different tasks can be jointly used to supervise the model training, so as to obtain a generic, versatile representation of point patterns that are informative about trees. Besides its proven abilities in terms of representation learning, the deep learning paradigm also offers end-to-end training from raw point cloud data, and thus a single, small set of hyper-parameters to tune the model to new forest and/or sensor characteristics. In our view, the segmentation problem is the critical part of the analysis. Once it has been solved with good accuracy, biophysical variables like tree height, DBH and crown attributes can be extracted with dedicated geometric algorithms, thus ensuring transparency and interpretability of the retrieval system. To validate the proposed approach, in the present study we:
@@
    Develop a deep network that ingests ALS-HD data and jointly performs semantic segmentation at plot/stand level (dominant canopy trees / intermediate and understory trees / low vegetation / ground), tree instance segmentation, and semantic component segmentation at tree level (stem / live branches / dead branches).
@@ 2.1 Dataset
Figure 1: Illustration of our segmentation and retrieval framework.
The FOR-Instance dataset (Puliti et al., 2023) is a recently published, machine learning-ready collection of 3D point clouds with manually annotated per-point semantic class labels and tree IDs. The dataset also has DBH values acquired by field measurement for selected trees. The point clouds were collected with drones and helicopters equipped with survey-grade laser scanners (Riegl VUX-1 UAV and Mini-VUX) at 67 plots across various geographic regions and forest types. For our study, we removed points with reference label outpoints, i.e., incomplete, partially observed trees along plot borders. This leaves us with five semantic categories: low vegetation, ground, stem points, live branches and dead branches, see Figure 1. Points on stems, live branches and dead branches are considered tree points, the other two categories are non-tree points. The data is split into 42 training plots, 14 validation plots and 11 test plots, in such a way that all five geographic regions are present in both the training and the test portions. [^1] The NIBIO2 region of FOR-Instance had to be excluded from the quantitative experiments, because it lacks DBH reference data, and because a significant number of understory trees are not annotated in the reference.
@@ 2.2 Deep learning framework for multiple segmentation tasks
Our forest inventory system consists of two parts, see Figure 1. First, a deep neural network labels the individual 3D points, simultaneously assigning semantic labels and instance IDs, using a shared feature extraction backbone. In a second step, learning-free geometric methods operate on the segmented data to retrieve a suite of forestry-related biophysical variables at the per-tree and per-stand levels.
2.2.1 Input data
@@
The input to the segmentation network is a point set $P\in\mathbb{R}^{N\times C}$, where $N$ is the number of points, and $C$ is the dimension of the per-point attributes. In our base setting, the attributes are only the $x$, $y$, $z$ coordinates, with the origin at the point cloud’s centre of gravity. If further useful information is available, e.g., additional sensor readings or forest-specific hand-crafted descriptors (Wang, 2020), it can be appended as additional attributes. For the presents study, we have explored the point-wise intensity, the return number and the scan angle rank, as given by the FOR-Instance dataset. Moreover, we have experimented with manually extracted geometric features derived from the eigenvalues of the local second moment matrix, including their sum, omnivariance, eigenentropy, anisotropy, planarity, linearity, surface variation, sphericity (Hackel et al., 2016) and verticality (Guinard and Landrieu, 2017). For efficiency, the entire point cloud is voxel grid subsampled to thin out overly dense regions and achieve a homogeneous (maximum) point density. The filter voxel size is set to 20$\times$20$\times$20 cm3, respectively 125 pts/m3.
@@
Training samples are drawn following a class-balanced random sampling scheme: 3D points are sampled with probability $P_{i}$ inversely proportional to the square root of the class frequency $N_{i}$ ($P_{i}\propto\sqrt{1/N_{i}}$). The sampled point defines the axis of a vertical cylinder of fixed radius, and the set of all points within the cylinder forms one training sample.
@@
In addition to the class-balanced sampling, we also experimented with three additional balancing strategies. Class weighting assigns weights inversely proportional to the square root of the per-class point count when computing the (semantic) segmentation loss, such that classes that are rare within a given sample receive higher weights. Height weighting assigns higher weights to low points (small $z$-coordinate), to reflect the uneven scan density along the vertical. The height values are normalized by their mean and the weight is inversely proportional to the logarithm of that height. Finally, for more balanced sampling across different regions, we test a region weighting scheme, where each geographic region is assigned a sampling probability, calculated from the region-wide class frequencies with the same inverse square root scheme as for class balancing.
@@
For each sampled cylinder region, various data augmentation techniques are applied during training: additive Gaussian random noise on the point coordinates (jittering), random rotations around the cylinder axis, random anisotropic scaling by factors $s\in[0.9,1.1]$, and random reflection along the $y$-axis. In addition, dropout is implemented for all input points: for every training sample, an unbiased coin flip (50% probability) decides whether to apply dropout; if yes, then 40% of the points in the sample are randomly discarded. That procedure aims to boost robustness against occlusions and missing points, especially for parts obscured by foliage and for the part near the ground, which tends to be sparser in ALS point clouds. Moreover, we apply random elastic deformation to enhance the representation of curved tree stems.

### 23 Feb 2024 v2

@@ Abstract
Detailed forest inventories are critical for sustainable and flexible management of forest resources, to conserve various ecosystem services. Modern airborne laser scanners deliver high-density point clouds with great potential for fine-scale forest inventory and analysis, but automatically partitioning those point clouds into meaningful entities like individual trees or tree components remains a challenge. The present study aims to fill this gap and introduces a deep learning framework Δ=Δ Δ=Δ Δ=Δ Δ=Δ , termed ForAINet, that is able to perform such a segmentation across diverse forest types and geographic regions. From the segmented data, we then derive relevant biophysical parameters of individual trees as well as stands. The system has been tested on FOR-Instance, a dataset of point clouds that have been acquired in five different countries using surveying drones. The segmentation back-end achieves over 85% F-score for individual trees, respectively over 73% mean IoU across five semantic categories: ground, low vegetation, stems, live branches and dead branches. Building on the segmentation results our pipeline then densely calculates biophysical features of each individual tree (height, crown diameter, crown volume, DBH, and location) and properties per stand (digital terrain model and stand density). Especially crown-related features are in most cases retrieved with high accuracy, whereas the estimates for DBH and location are less reliable, due to the airborne scanning setup.
@@ 1 Introduction
Collecting and maintaining forest inventories was one of the first applications to adopt airborne laser scanning (ALS) as an operational tool, at local (Næsset et al., 2004), regional and national scales (Kangas et al., 2018). The two main approaches to extract forest inventory data from Light Detection and Ranging (LiDAR) point clouds are the area-based approach (ABA, Næsset, 2002) and individual tree detection (ITD, Hyyppa et al., 2001). ABA operates by estimating statistical indicators from the points within a given area (e.g., a grid cell). In contrast, ITD (also referred to as single tree inventory) relies on the detection and segmentation of individual trees in the point cloud (Persson et al., 2002) Δ=Δ Δ=Δ Δ=Δ Δ=Δ , and consequently requires data at higher resolution (Coomes et al., 2017). Several methods have been developed for ITD, whose accuracies differ depending on the forest conditions and the point density (Vauhkonen et al., 2011; Kaartinen et al., 2012). Most ITD algorithms operate on a rasterised canopy height model (CHM) derived from the highest ALS returns, representing the canopy top layer. Consequently, they are limited to the dominant canopy trees and tend to miss an important share of intermediate and understory trees. Despite the large body of research on the topic, the operational uptake of ITD inventories has so far been limited, due to (1) the bias introduced by the relatively low detection rates, caused by the omission of non-dominant trees, and (2) a lack of transferability across different forest conditions and datasets.
@@
Δ=Δ Δ=Δ Δ=Δ Δ=Δ Traditionally, low-resolution ALS-based inventories have been based on rather sparse point clouds (0.5 – 5 pts/m2) for ABA, and slightly denser for ITD (10 – 20 pts/m2). With the rapid development of sensor technology, very high density ALS (ALS-HD) has become available Δ=Δ Δ=Δ Δ=Δ Δ=Δ , with point cloud densities ranging from 500 up to 10,000 pts/m2, offering new opportunities (Kellner et al., 2019; Puliti et al., 2022). Δ=Δ Δ=Δ Δ=Δ Δ=Δ While also collected with airborne platforms such as drones (as in the present study) or helicopters (Hyyppä et al., 2022; Persson et al., 2022), ALS-HD data differ from conventional ALS data in that they are acquired from lower altitude (and at lower flight speed), and usually with larger overlap between adjacent flight strips. Despite covering smaller geographical regions (e.g., forest stands), ALS-HD represents an emerging alternative data source for ITD inventories due to their (for airborne data) unprecedented level of detail, closing the gap between large-area ALS and high-resolution terrestrial laser scanning (TLS).
@@
The increased amount of 3D structural detail, with distinctively visible individual trees throughout the vertical canopy profile, has sparked renewed interest in the pursuit of single tree inventories, and possibly even completely airborne inventories based on dense retrievals of the variables of interest from the point clouds (Liang et al., 2019; Puliti et al., 2020). In addition to the measurement of common forestry variables like the diameter at breast height (DBH, Kuželka et al., 2020) and tree biomass (Brede et al., 2019), recent research has highlighted the potential to derive a suite of useful tree analytics and insights from per-tree ALS-HD data, such as species (Hakula et al., 2023), stem curve (Hyyppä et al., 2022), growth trajectory (Puliti et al., 2022), and breeding-related traits (du Toit et al., 2023). The promise of high-resolution scans for ITD is evident, but there is a bottleneck: to measure the properties of each tree, one first needs a sufficiently accurate and transferable method to segment the raw scan data into individual trees.
@@
Parallel to the rise of high-density ALS, we have seen an explosion of interest in deep machine learning algorithms for processing 3D point clouds in general, and forest scans in particular (Table 1). Over the last decade, the superior capabilities of deep neural networks when it comes to representing and analysing structured data have brought about transformative changes across a range of science and engineering disciplines (Wang et al., 2023a). Deep learning also holds great promise for tree segmentation, and indeed there has been a recent trend to adopt it for the ongoing efforts towards automated forest inventories (Xi et al., 2018; Chen et al., 2021b; Krisanski et al., 2021b; Sun et al., 2022; Chang et al., 2022; Wang and Bryson, 2023; Zhang et al., 2023; Kim et al., 2023; Jiang et al., 2023; Wielgosz et al., 2023; Straker et al., 2023). Δ=Δ Δ=Δ Δ=Δ Δ=Δ The review in Table 1 indicates that most methods either apply 2D CNNs on CHM or use PointNet++ (Qi et al., 2017) to directly process point clouds. Networks that operate on the CHM share the limitations of traditional 2.5D ITD algorithms, i.e., they disregard understory trees. The seminal PointNet++ architecture does consider all scan points, including those below the main canopy. But new network designs, particularly sparse 3D convolutional networks (Choy et al., 2019), offer higher accuracy as well as computational advantages, making them more suitable for high-density point clouds.
@@
Δ=Δ Δ=Δ Δ=Δ Δ=Δ Data augmentation synthetically enhances the diversity of the training data and prevents overfitting, thus improving model accuracy without the need to collect additional field data (Chen et al., 2021b; Zhang et al., 2023). The methods in Table 1 usually employ basic data augmentation techniques like rotation, mirroring and coordinate jittering (Chen et al., 2021b; Krisanski et al., 2021b; Kaijaluoto et al., 2022). Yet, we are not aware of data augmentation strategies that take into account the specific characteristics of LiDAR point clouds captured in forests.
Another significant barrier to the advancement of automated forest inventory tools has been the lack of large reference datasets. Table 1 highlights the limited availability of open data that, moreover, often provide only semantic category labels but not tree instance labels. The scarcity of labeled data for supervised learning presents a bottleneck for the use of deep learning. The emergence of comprehensive labeled datasets like FOR-Instance is an important prerequisite to enable 3D deep learning in support of forest inventories.
In the present work, we consolidate data-driven segmentation and structuring of forest point clouds into a single, integrated deep learning framework. We argue that a key ingredient for an efficient, automatic airborne forest inventory and management system is a comprehensive segmentation engine Δ=Δ Δ=Δ Δ=Δ Δ=Δ , tuned specifically to the data characteristics of forest scans. In this way, reference data for different tasks can be jointly used to supervise the model training, so as to obtain a generic, versatile representation of point patterns that are informative about trees. Besides its proven abilities in terms of representation learning, the deep learning paradigm also offers end-to-end training from raw point cloud data, and thus a single, small set of hyper-parameters to tune the model to new forest and/or sensor characteristics. In our view, the segmentation problem is the critical part of the analysis. Once it has been solved with good accuracy, biophysical variables like tree height, DBH and crown attributes can be extracted with dedicated geometric algorithms, thus ensuring transparency and interpretability of the retrieval system. To validate the proposed approach, in the present study we:
@@
    Develop a deep network Δ=Δ Δ=Δ Δ=Δ Δ=Δ , ForAINet, that ingests ALS-HD data and jointly performs semantic segmentation at plot/stand level (dominant canopy trees / intermediate and understory trees / low vegetation / ground), tree instance segmentation, and semantic component segmentation at tree level (stem / live branches / dead branches).
@@ 2.1 Dataset
Figure 1: Illustration of our segmentation and retrieval framework. Δ=Δ Δ=Δ Δ=Δ Δ=Δ It operates in two steps: The first step segments points into semantic categories as well as individual trees, for details see Figure 2 and Section 2.2. The second step retrieves tree parameters and stand structure from the segmentation results, see Section 2.3.
Table 2: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Characteristics of the FOR-Instance dataset in different geographic regions.
[…]
Train
Eval
Test
CULS
Coniferous dominated temperate forest
Pinus silvestris
Riegl VUX-1 UAV
1
1
1
47
2585
Czech Republic
NIBIO
Coniferous dominated boreal forest
Picea abies (dominated)
Pinus Sylvestris (few)
Betula pendula (few)
Riegl MiniVUX-1 UAV
8
6
6
575
9529
Norway
RMIT
Native dry sclerophyll eucalypt forest
Eucalyptus pulchella
Riegl MiniVUX-1 UAV
1
0
1
223
498
Australia
SCION
Non-native pure coniferous temperate forest
Pinus radiata
Riegl MiniVUX-1 UAV
2
1
2
135
4576
New Zealand
TUWIEN
Deciduous dominated alluvial forest
Deciduous species
Riegl VUX-1 UAV
1
0
1
150
1717
Austria
The FOR-Instance dataset (Puliti et al., 2023) is a recently published, machine learning-ready collection of 3D point clouds with manually annotated per-point semantic class labels and tree IDs. The dataset also has DBH values acquired by field measurement for selected trees. The point clouds were collected with drones and helicopters equipped with survey-grade laser scanners (Riegl VUX-1 UAV and Mini-VUX) at 67 plots across various geographic regions and forest types. Δ=Δ Δ=Δ Δ=Δ Δ=Δ For our study, points with reference label outpoints, corresponding to incomplete or partially observed trees along plot borders, were removed due to the absence of instance labels for them. This leaves us with five semantic categories: low vegetation, ground, stem points, live branches and dead branches, see Figure 1. Points on stems, live branches and dead branches are considered tree points, the other two categories are non-tree points. The data is split into 42 training plots, 14 validation plots and 11 test plots, in such a way that all five geographic regions are present in both the training and the test portions. [^1] The NIBIO2 region of FOR-Instance had to be excluded from the quantitative experiments, because it lacks DBH reference data, and because a significant number of understory trees are Δ=Δ Δ=Δ Δ=Δ Δ=Δ not annotated in the reference. Δ=Δ Δ=Δ Δ=Δ Δ=Δ Table 2 lists important characteristics of the FOR-Instance dataset, separately per geographic region. For further information about the data we refer the reader to Puliti et al. (2023).
@@ 2.2 Deep learning framework for multiple segmentation tasks
Our forest inventory system consists of two parts, see Figure 1. First, a deep neural network labels the individual 3D points, simultaneously assigning semantic labels and instance IDs, using a shared feature extraction backbone Δ=Δ Δ=Δ Δ=Δ Δ=Δ (Figure 2). Δ=Δ Δ=Δ Δ=Δ Δ=Δ For brevity, we call that component ForAINet, short for Forest Automatic Inventory Neural Network. In a second step, learning-free geometric methods operate on the segmented data to retrieve a suite of forestry-related biophysical variables at the per-tree and per-stand levels.
Figure 2: Δ=Δ Δ=Δ Δ=Δ Δ=Δ Illustration of the point cloud segmentation pipeline.
2.2.1 Δ=Δ Δ=Δ Δ=Δ Δ=Δ Data augmentation and balancing strategies
@@
The input to the segmentation network is a point set $P\in\mathbb{R}^{N\times C}$, where $N$ is the number of points, and $C$ is the dimension of the per-point attributes. In our base setting, the attributes are only the $x$, $y$, $z$ coordinates, with the origin at the point cloud’s centre of gravity. If further useful information is available, e.g., additional sensor readings or forest-specific hand-crafted descriptors (Wang, 2020), it can be appended as additional attributes. For the presents study, we have explored the point-wise intensity, the return number and the scan angle rank, as given by the FOR-Instance dataset. Moreover, we have experimented with Δ=Δ Δ=Δ Δ=Δ Δ=Δ traditional, hand-crafted geometric features derived from the eigenvalues of the local second moment matrix, including their sum, omnivariance, eigenentropy, anisotropy, planarity, linearity, surface variation, sphericity (Hackel et al., 2016) and verticality (Guinard and Landrieu, 2017) Δ=Δ Δ=Δ Δ=Δ Δ=Δ , c.f. Section 2.5. For efficiency, the entire point cloud is voxel grid subsampled Δ=Δ Δ=Δ Δ=Δ Δ=Δ to a single, randomly chosen point per voxel, so as to thin out overly dense regions and achieve a homogeneous (maximum) point density. The filter voxel size is set to 20$\times$20$\times$20 cm3, respectively 125 pts/m3.
@@
Training samples are drawn following a class-balanced random sampling scheme: 3D points are sampled with probability $P_{i}$ inversely proportional to the square root of the class frequency $N_{i}$ ($P_{i}\propto\sqrt{1/N_{i}}$) Δ=Δ Δ=Δ Δ=Δ Δ=Δ , where the subscript $i$ indicates the five different semantic classes. The sampled point defines the axis of a vertical cylinder of fixed radius, and the set of all points within the cylinder forms one training sample. Δ=Δ Δ=Δ Δ=Δ Δ=Δ In our implementation the radius is set to 8 m, enough to ensure almost every tree in the FOR-Instance data is contained in a single cylinder (for other biomes this value may have to be adapted).
@@
In addition to the class-balanced sampling, we also experimented with three additional balancing strategies Δ=Δ Δ=Δ Δ=Δ Δ=Δ , in order to determine and adopt the most effective strategy. Class weighting assigns weights inversely proportional to the square root of the per-class point count when computing the (semantic) segmentation loss, such that classes that are rare within a given sample receive higher weights. Height weighting assigns higher weights to low points (small $z$-coordinate), to reflect the uneven scan density along the vertical. The height values are normalized by their mean and the weight is inversely proportional to the logarithm of that height. Finally, for more balanced sampling across different regions, we test a region weighting scheme, where each geographic region is assigned a sampling probability, calculated from the region-wide class frequencies with the same inverse square root scheme as for class balancing.
@@
For each sampled cylinder region, various data augmentation techniques are applied during training: additive Gaussian random noise on the point coordinates (jittering), random rotations around the cylinder axis, random anisotropic scaling by factors $s\in[0.9,1.1]$, and random reflection along the $y$-axis. In addition, Δ=Δ Δ=Δ Δ=Δ Δ=Δ a subsampling strategy is implemented for all input points: for every training sample, an unbiased coin flip (50% probability) decides whether to apply Δ=Δ Δ=Δ Δ=Δ Δ=Δ subsampling; if yes, then 40% of the points in the sample are randomly discarded. That procedure aims to boost robustness against occlusions and missing points, especially for parts obscured by foliage and for the part near the ground, which tends to be sparser in ALS point clouds. Moreover, Δ=Δ Δ=Δ Δ=Δ Δ=Δ elastic deformation is applied, with the aim to enhance the recognition of curved tree stems. To that end, a field of Gaussian random deformations with magnitude $\in[0.4,1.6]\,$m is sampled on a regular grid of granularity $\in[0.2,0.8]\,$m. The field is then smoothed to ensure spatial coherence and applied to the point coordinates via trilinear interpolation, causing local bending of previously straight structures such as stems.